### v2

# Pilot-waves and copilot-particles: A novel approach to objective collapse

Abstract
We propose an extension of Schrödinger’s equation that incorporates the macroscopic measurement-induced wavefunction collapse phenomenon. Our approach relies on a hybrid between Bohm-de Broglie pilot-wave and objective collapse theories. The Bohmian particle is guided by the wavefunction and, conversely, the wavefunction gradually localizes towards the particle’s position. As long as the particle can visit any state, as in a typical microscopic system, the localization effect does not favor any particular quantum state and, on average, the usual Schrödinger-like time evolution results. However, when the wavefunction develops spatially well-separated lobes, as would happen during a macroscopic measurement, the Bohmian particle can remain trapped in one lobe, and the wavefunction eventually localizes there. The end result, in macroscopic systems, is a wavefunction collapse that is consistent with Born’s rule. We illustrate the theory with a simple double-slit experiment simulation.
Over a century after the birth of quantum mechanics, the search for a compelling understanding of wavefunction “collapse” in quantum mechanics is still ongoing [1, 2, 3, 4]. Existing proposals include decoherence [5, 6], pilot-waves or Bohmian mechanics [7, 8, 9], many-world interpretations [10], Bayesian interpretations [11], objective collapse [12], gravitational effects [13, 14], and other emerging ideas [15, 16, 17, 18]. These efforts have resulted in significant conceptual progress but no single approach has been universally accepted as a solution to the measurement problem in quantum mechanics [19].
As a step towards this goal, we propose a hybrid between Bohm-de Broglie pilot-wave [9] and objective collapse theories [12] that incorporate a form of wavefunction collapse consistent with Born’s rule. Our approach aims to address some of the conceptual questions associated with either approaches: How do observers perceive the Bohmian particle when a physical system’s behavior only depends on its wavefunction? What determine when and where the wavefunction objectively collapses?
In analogy with pilot-wave theories, we posit that the global wavefunction drives a Bohmian “particle”. Our approach however postulates a feedback in the other direction as well: The wavefunction gradually localizes towards the particle’s position. In the limit in which the particle speed is high relative to the rate of localization, these modifications to the Schrödinger equation have little effect on the fairly localized wavefunction of microscopic systems. But when the wavefunction develops multiple macroscopically distinct regions of high probability (hereafter, “lobes”), the particle may become trapped in one of these regions and the wavefunction starts to localize towards only one of them. Since the probability of “selecting” one lobe is equal to the probability of the particle lying in this lobe at the moment the different parts of the wavefunction lose connectivity, Born’s rule arises from this process as well. Obtaining this trapping behavior in the macroscopic limit demands the use of different equation of motion than those of conventional pilot-wave theory.
The theory is more naturally presented in the context of the Schrödinger equation for continuous degrees of freedom (see End Matter for a treatment of discrete states). Below, $\psi\left(x,t\right)$ is a properly (anti-)symmetrized wavefunction in position representation, with $t$ being time and $x$ being a $3N$-dimensional vector for a system of $N$ particles. (We consider the familiar context of “first quantization”, but there is no fundamental problem associated with applying it to boson fields within quantum field theory [20], in which case $x$ would represent a specific value of a classical field.) Like most theories with either a pilot-wave or an objective collapse flavor, our theory is non-relativistic. Our proposed modified Schrödinger equation takes the form:
$\frac{d\psi\left(x,t\right)}{dt}=\frac{1}{\mathbf{i}\hbar}H\psi\left(x,t\right)-\kappa\psi\left(x,t\right)+\kappa\frac{\bar{\phi}_{r\left(t\right)}\left(x\right)\psi\left(x,t\right)}{\left|\psi\left(r\left(t\right),t\right)\right|^{2}}$ (1)
where the Hamiltonian has the usual form $H=-\frac{\hbar^{2}}{2}m^{-1}\nabla^{2}+V\left(x\right)$ for a given imposed potential $V\left(x\right)$ and a given diagonal matrix of masses $m$. The first term of the right-hand side on (1) is standard while the remaining two implement the localization of the wavefunction towards the particle located at $r\left(t\right)$ (in $\mathbb{R}^{3N}$) at time $t$. The second term introduces a tendency for exponential decay of the magnitude of the wavefunction everywhere, without affecting its phase. The parameter $\kappa$ controls the rate of this process. In the last term, $\bar{\phi}_{r}\left(x\right)$ is smooth function that only takes on a large value near the position $r$ or the Bohmian particle: $\bar{\phi}_{r}\left(x\right)=\frac{1}{\#\Pi}\sum_{\pi\in\Pi}\sigma^{-3N}\phi\left((x-\pi r)/\sigma\right)$, where $\phi\left(u\right)$ is a localized function normalized to integrate to $1$ (e.g. a standard Gaussian), $\sigma$ controls the spatial extent of the localization and $\Pi$ is the set of coordinate permutations among identical particles. This symmetrization device ensures that the localization effect does not destroy the symmetry or anti-symmetry of the wavefunction. Note that, since neither of the last two terms affect the phase of the wavefunction, both Fermions and Bosons can be handled in the same way: the sign alternation needed for Fermions is already embedded in the wavefunction and there is no need for sign alternation in the $\bar{\phi}_{r}\left(x\right)$ term. We also implement the localization through a smooth function $\phi\left(u\right)$ instead of a, say, a delta function, to avoid generating high-frequency noise in the wavefunction that, if present, would likely have been experimentally detected [3]. The net effect of the last term is an increase in wavefunction magnitude near the particle. The normalization by $\left|\psi\left(r,t\right)\right|^{2}$ ensures that, on average, the last two terms cancel out, as we will demonstrate shortly.
There remains to specify to equation of motion for $r$. We depart from the standard Bohm-de Broglie prescription and instead assume that the particle experiences a fictitious potential $-\tau\ln\left|\psi\left(r,t\right)\right|^{2}$, where $\tau$ is a constant, and otherwise obeys a classical equation of motion:
$\frac{d^{2}r}{dt^{2}}=\mu^{-1}\tau\nabla_{r}\ln\left|\psi\left(r,t-\Delta t\right)\right|^{2},$ (2)
where $\mu$ is a $3N\times 3N$ diagonal matrix of fictitious masses. We allow for a time delay $\Delta t$ in the effect of the wavefunction (having $\left|\psi\left(r,t-\Delta t\right)\right|$ instead of the instantaneous effect $\left|\psi\left(r,t\right)\right|$) to avoid a self-interaction artifact (to be discussed in more detail below). Standard statistical mechanics arguments [21] imply that, in equilibrium at a fictitious temperature $\tau$ (expressed in energy units), the position $r$ will be distributed as $\exp\left(-\left(-\tau\ln\left|\psi\left(r,t\right)\right|^{2}\right)/\tau\right)=\left|\psi\left(r,t\right)\right|^{2}$, as desired. This presupposes that the velocities are initialized (in the distant past) consistently with the fictitious temperature $\tau$, e.g. $\frac{dr_{k}}{dt}\sim\operatorname{Normal}\left(0,\mu_{k}^{-1}\tau\right)$ for $k=1,\ldots,3N$. This argument also assumes that the particle equilibrates rapidly relative to the time scale of changes in the wave function magnitude $\left|\psi\left(r,t\right)\right|$. Note that all “fictitious” quantities are not tied to any real observable properties of the system: $\mu$ is not necessarily related to the actual particles’ masses and $\tau$ may not be the real temperature. In fact, these quantities enter all equations as the ratio $\mu^{-1}\tau$, which we leave as an adjustable parameter in the theory. It is interesting that the logarithm in Equation (2) coincidentally accomplishes two tasks: separating out the magnitude from the phase (since $\ln\left|\psi\left(r,t\right)\right|^{2}=2\operatorname{Re}\ln\psi\left(r,t\right))$ and placing the magnitude on a logarithmic scale, as needed to cancel out the Boltzmann factor.
Figure 1: (Color online) Microscopic vs. macroscopic behaviors. Upper left: For a spatially localized wavefunction, the Bohmian particle (represented by a grey spheres trail) can access all regions where the wavefunction is nonnegligible (here represented by fictitious potential surface proportional to $-\ln|\psi(r,t)|^{2}$). Middle: As a superposition of states starts to involve spatially well-separated states, the particle remains trapped in one of the “lobes” of the wavefunction. Lower right: Loss of ergodicity causes wavefunction localization around one of the macroscopically distinct outcomes
The reason we deviate from standard Bohm-de Broglie dynamics are twofold. First, the particles tend to not get trapped very well under the usual Bohm-de Broglie dynamics: As the particle approaches a region of low probability density, it typically speeds up (which specifically ensures that the particle has low probability of being found there). As a result, the particle can easily escape any isolated lobe in the wavefunction. Second, the Bohmian particle does not thermalize very quickly. For instance, in a wave packet, the particle typically follows along smoothly while remaining at the same location relative to the center of the wave packet, instead of visiting all positions of high probability. Our alternative dynamics avoids both of these problems.
We are now ready to show that, under conditions ensuring thermalization, our proposal essentially reduces to the conventional Schrödinger equation. Indeed, computing the expected value of the last term of (1) with respect to the random position $r$ with instantaneous probability density $\left|\psi\left(r,t\right)\right|^{2}$, we have:
[…]
In contrast, as illustrated in Figure 1, if the wavefunction develops spatially well-separated lobes and the particle is trapped in only one of the them, then the last term remains essentially zero for the lobes that are no longer visited and, there, only the decaying term $-\kappa\psi\left(x,t\right)$ is active, resulting in an exponential decay of the wavefunction. Within the lobe trapping the particle, the localization term of (1) is larger than average, since the particle visit points there in more often than would be indicated by the wavefunction magnitude. The end result is that the localization terms there exceeds the decaying term on average and the wavefunction gradually concentrates onto that lobe at the expense of the others. The loss of ergodicity of the Bohmian particle is the source of the wavefunction collapse within this model. This helps explain single outcomes in macroscopic measurements, since they arguably lead to wavefunctions with macroscopically disconnected lobes, such as a dial (or “pointer”) being in one place or another or a photon being emitted in different directions, etc. The End Matter extends this reasoning to the measurement of a subsystem.
Similarly, another mechanism that could cause a loss of ergodicity is the process of copying a microscopic state onto a large number of particles (within the limits of the no-cloning theorem). This operation typically results in increasingly deeper basins in the fictitious potential $-\tau\ln\left|\psi\left(r,t\right)\right|^{2}$ that increase the likelihood of the particle being trapped. To see this, consider a particle in a superposition of two states $\alpha\left|1\right\rangle+\beta\left|2\right\rangle$ (with $\left|\alpha\right|^{2}+\left|\beta\right|^{2}=1$) that is starting to interact with a measurement apparatus made of $N-1$ distinguishable particles, initially in a “no measurement” state $\left|0\right\rangle\otimes\cdots\otimes\left|0\right\rangle$. After a suitable unitary evolution, the joint system transitions to the state $\alpha\left|1\right\rangle\otimes\left|1\right\rangle\otimes\cdots\otimes\left|1\right\rangle+\beta\left|2\right\rangle\otimes\left|2\right\rangle\otimes\cdots\otimes\left|2\right\rangle$. Now assume that, for particle $i$, the states $\left|1\right\rangle$ and $\left|2\right\rangle$ are, respectively, associated with single-particle wavefunctions in position representation $\psi_{1}\left(x_{i}\right)=\varphi\left(x_{i}-a\right)$ and $\psi_{2}\left(x_{i}\right)=\varphi\left(x_{i}+a\right)$ for some macroscopic value $a\in\mathbb{R}^{3}$ and centrosymmetric unimodal function $\varphi\left(x_{i}\right)$ (with $x_{i}\in\mathbb{R}^{3}$). The overall wavefunction is then
$\psi\left(x\right)=\alpha\prod\nolimits_{i=1}^{N}\varphi\left(x_{i}-a\right)+\beta\prod\nolimits_{i=1}^{N}\varphi\left(x_{i}+a\right).$
To estimate the escape rate out of one well, we calculate the fictitious energy barrier $\Delta\epsilon$, defined as the difference between the particle’s fictitious energy at the saddle point at $x_{i}=0$ (for all $i$) and at the bottom of one of the wells (say, at $x_{i}=a$, for all $i$): $\Delta\epsilon\equiv C+N2\tau\ln\left(\left|\varphi\left(0\right)\right|/\left|\varphi\left(a\right)\right|\right)$, where we assumed $\varphi\left(2a\right)\approx 0$ and $C$ is a constant independent of $N$. Since the barrier increases linearly with $N$, the transition rate between the two wells is decreasing exponentially with the number $N$ of particles involved in the measurement.
Figure 2 illustrates our proposed time evolution of the wavefunction in a classic setting: the double-slit experiment. The simulation shows both the initial wave-like interference behavior and the eventual collapse to a more specific outcome as the wave packet heads towards a detector. The Supplementary materials include an animation of the same numerical experiment, as well as the simulation code itself. The End Matter also includes a treatment of a standard EPR-type experiment [22].
[…]
The proposed theory improves upon the Bohm-de Broglie pilot-wave approach along a few dimensions. First, it avoids the awkwardly distinct treatment of (i) the physical system being described by the wavefunction and (ii) the observer only perceiving the position of the Bohmian particle. Having a feedback mechanism through which the particle affects the wavefunction solves this problem and also leads to a model of wavefunction collapse. In our approach, both the wavefunction and the particle jointly describe the behavior of both the physical system and the observer. The definite state of the particle is what drives the wavefunction localization during the measurement process. A conceptually appealing feature of the theory is that avoids the often criticized exponential growth of alternative realities in which the Bohmian particle does not reside and that seems entirely superfluous to describe the actual state of the world. In our approach, the wavefunction associated with these alternative realities simply die out exponentially fast.
The proposed theory also improves some aspects of objective collapse theories, in which the position of the collapse is simply picked at random based on the wavefunction and a random noise field. On a conceptual level, our approach provides a rationale for the origin of this randomness: The random location of the collapse is simply determined by the random location of the particle. The type of collapse we propose may also be more amenable to an eventual relativistic treatment, as we shall discuss further below. On a more empirical level, a problem with objective collapse theories is the abruptness of the collapse process, which is associated with a high-frequency signature which has yet to be found experimentally [3]. In our approach, the collapse is a smoother process more consistent with experimental tests.
The idea that the wavefunction is attracted to particles can be found in earlier work [18], but our approach differs in substantial ways. We consider collapse towards a point in $\mathbb{R}^{3N}$ instead of a towards a single-particle density. Keeping $3N$-dimensional information helps maintain EPR-type correlations between distant measurements. Also, unlike [18], we deviate from standard Bohmian dynamics to ensure particle trapping.
The proposed theory is compatible with the standard decoherence picture (see End Matter for a density matrix formulation): it does not affect the basic idea that interactions with the environment causes a decay of the off-diagonal terms of a subsystem’s density matrix. But once the resulting mixture of outcomes involves macroscopically distinct states, our equation of motion predicts localization towards one of these macroscopic states. Hence, it provides an objective “selection rule” that is absent under decoherence alone and avoids the need for a many-world interpretation. The decoherence concept makes it unnecessary for the wave function to collapse rapidly in our framework. The conversion from a superposition of states to a statistical mixture takes place over the same time span under our theory as under decoherence alone. The exponential decay of all but one branch of the statistical mixture, however, can take much longer: any decay rate faster than human perception (i.e., milliseconds) would be sufficient to give human observers the experience of a single measurement outcome.
The proposed theory does exhibit some undesirable features (see End Matter). Most obviously, the proposed modified Schrödinger equation is slightly nonlinear, whereas linearity of quantum mechanics has been verified experimentally to a great precision [23]. This places a tight upper bound on the parameter $\kappa$, thus also placing a bound on the speed at which wavefunction collapse can occur. The fact that the nonlinearity must be small mitigates other known issues associated with nonlinearity [24].
While the current treatment is non-relativistic, nothing seems to forbid a relativistic extension. The collapse phenomenon involves local operations: the decay term $-\kappa\psi\left(x,t\right)$ is purely local while the localization terms only involves instantaneous effects over short distances (as governed by the function $\phi_{r}\left(x\right)$) and ensuring that this effect propagates at a finite speed entails no significant changes in the theory. Since relativistic versions of pilot-wave [25] and objective collapse theories [26] have been developed with some success, similar ideas could carry over in the present setup.
In our theory, the time delay $\Delta t$ should satisfy $\Delta t\gg\sigma/\sqrt{\tau/\mu}$ (with scalar mass $\mu$ for simplicity) to avoid a strong instantaneous self-interaction. In this fashion, the wavefunction localization only starts to have an effect on the particle when it has already moved by $\Delta t$ times its average velocity $\sqrt{\tau/\mu}$. If this distance exceeds $\sigma$, the particle will not experience the “indentation” it has just created in the wavefunction. Absent of this delay, the particle would have a tendency to gradually lose kinetic energy. This is not the only way to address the self-interaction problem: one could add a negative drag term in the equation of motion for $r$ instead, or place the particle in weak interaction with an outside reservoir that constantly replenishes the lost energy to maintain an average kinetic energy $\tau$. We favor the time delay approach for its simplicity, but it would be difficult to discern these alternatives experimentally.
To summarize, the key new aspects of the theory include: (i) a mutual feedback between the Bohmian particle and the wavefunction; (ii) a different proposed equation of motion for the Bohmian particle; (iii) a mechanism based on the loss of ergodicity (or trapping) of the Bohmian particle to explain to wavefunction collapse phenomena; (iv) the idea that the collapse mechanism can actually be gradual and local. The fact the proposed evolution equations are slightly nonlinear, could enable tests of the proposed theory, as could situations where the Bohmian particle would oscillate in resonance with oscillations in the wavefunction. Confirmation would shed light on some of the conceptual foundation of quantum mechanics and could cast doubt over the feasibility of large-scale quantum computers.