### v3

# Pilot-waves and copilot-particles: A nonstochastic approach to objective wavefunction collapse

Abstract
We seek an extension to Schrödinger’s equation that incorporates the macroscopic measurement-induced wavefunction collapse phenomenon. We find that a suitable hybrid between two leading approaches, the Bohm-de Broglie pilot-wave and objective collapse theories, accomplishes this goal in a way that is consistent with Born’s rule. Our theory posits that the Bohmian particle is guided by the wavefunction and, conversely, the wavefunction gradually localizes towards the particle’s position. As long as the particle can visit any state, as in a typical microscopic system, the localization effect does not favor any particular quantum state and, on average, the usual Schrödinger-like time evolution results. However, when the wavefunction develops spatially well-separated lobes, as would happen during a macroscopic measurement, the Bohmian particle can remain trapped in one lobe, which causes the wavefunction to eventually localizes. This proposed loss of ergodicity mechanism recasts one of the foundational postulate of quantum mechanics as a emergent feature and has important implications regarding the feasibility of large-scale quantum computing.
[…]
Measurement postulate, Bohmian mechanics, Nonlinear wave equation
1 Introduction
Over a century after the birth of quantum mechanics, the search for a compelling understanding of wavefunction “collapse” in quantum mechanics is still ongoing [1, 2, 3, 4]. Existing proposals include decoherence [5, 6], pilot-waves or Bohmian mechanics [7, 8, 9], many-world interpretations [10], Bayesian interpretations [11], objective collapse [12], gravitational effects [13, 14], and other emerging ideas [15, 16, 17, 18]. These efforts have resulted in significant conceptual progress but no single approach has been universally accepted as a solution to the measurement problem in quantum mechanics [19].
As a step towards this goal, we propose a hybrid between Bohm-de Broglie pilot-wave [9] and objective collapse theories [12] that incorporate a form of wavefunction collapse consistent with Born’s rule. Our approach aims to address some of the conceptual questions associated with either approaches: How do observers perceive the Bohmian particle when a physical system’s behavior only depends on its wavefunction? What determine when and where the wavefunction objectively collapses?
In analogy with pilot-wave theories, we posit that the global wavefunction drives a Bohmian “particle”. Our approach however postulates a feedback in the other direction as well: The wavefunction gradually localizes towards the particle’s position. In the limit in which the particle speed is high relative to the rate of localization, these modifications to the Schrödinger equation have little effect on the fairly localized wavefunction of microscopic systems. But when the wavefunction develops multiple macroscopically distinct regions of high probability (hereafter, “lobes”), the particle may become trapped in one of these regions and the wavefunction starts to localize towards only one of them. Since the probability of “selecting” one lobe is equal to the probability of the particle lying in this lobe at the moment the different parts of the wavefunction lose connectivity, Born’s rule arises from this process as well. Obtaining this trapping behavior in the macroscopic limit demands the use of different equation of motion than those of conventional pilot-wave theory.
2 Theory
The theory is more naturally presented in the context of the Schrödinger equation for continuous degrees of freedom (see Appendix E for a treatment of discrete states). Below, $\psi\left(x,t\right)$ is a properly (anti-)symmetrized wavefunction in position representation, with $t$ being time and $x$ being a $3N$-dimensional vector for a system of $N$ particles. (We consider the familiar context of “first quantization”, but there is no fundamental problem associated with applying it to boson fields within quantum field theory [20], in which case $x$ would represent a specific value of a classical field.) Like most theories with either a pilot-wave or an objective collapse flavor, our theory is non-relativistic. Our proposed modified Schrödinger equation takes the form:
$\frac{d\psi\left(x,t\right)}{dt}=\frac{1}{\mathbf{i}\hbar}H\psi\left(x,t\right)-\kappa\psi\left(x,t\right)+\kappa\frac{\bar{\phi}_{r\left(t\right)}\left(x\right)\psi\left(x,t\right)}{\left|\psi\left(r\left(t\right),t\right)\right|^{2}}$ (1)
where the Hamiltonian has the usual form $H=-\frac{\hbar^{2}}{2}m^{-1}\nabla^{2}+V\left(x\right)$ for a given imposed potential $V\left(x\right)$ and a given diagonal matrix of masses $m$. The first term of the right-hand side on (1) is standard while the remaining two implement the localization of the wavefunction towards the particle located at $r\left(t\right)$ (in $\mathbb{R}^{3N}$) at time $t$. The second term introduces a tendency for exponential decay of the magnitude of the wavefunction everywhere, without affecting its phase. The parameter $\kappa$ controls the rate of this process. In the last term, $\bar{\phi}_{r}\left(x\right)$ is smooth function that only takes on a large value near the position $r$ or the Bohmian particle: $\bar{\phi}_{r}\left(x\right)=\frac{1}{\#\Pi}\sum_{\pi\in\Pi}\sigma^{-3N}\phi\left((x-\pi r)/\sigma\right)$, where $\phi\left(u\right)$ is a localized function normalized to integrate to $1$ (e.g. a standard Gaussian), $\sigma$ controls the spatial extent of the localization and $\Pi$ is the set of coordinate permutations among identical particles. This symmetrization device ensures that the localization effect does not destroy the symmetry or anti-symmetry of the wavefunction. Note that, since neither of the last two terms affect the phase of the wavefunction, both Fermions and Bosons can be handled in the same way: the sign alternation needed for Fermions is already embedded in the wavefunction and there is no need for sign alternation in the $\bar{\phi}_{r}\left(x\right)$ term. We also implement the localization through a smooth function $\phi\left(u\right)$ instead of a, say, a delta function, to avoid generating high-frequency noise in the wavefunction that, if present, would likely have been experimentally detected [3]. The net effect of the last term is an increase in wavefunction magnitude near the particle. The normalization by $\left|\psi\left(r,t\right)\right|^{2}$ ensures that, on average, the last two terms cancel out, as we will demonstrate shortly.
There remains to specify to equation of motion for $r$. We depart from the standard Bohm-de Broglie prescription and instead assume that the particle experiences a fictitious potential $\tau\ln\left|\psi\left(r,t\right)\right|^{2}$, where $\tau$ is a constant, and otherwise obeys a classical equation of motion:
$\frac{d^{2}r}{dt^{2}}=\mu^{-1}\tau\nabla_{r}\ln\left|\psi\left(r,t-\Delta t\right)\right|^{2},$ (2)
where $\mu$ is a $3N\times 3N$ diagonal matrix of fictitious masses. We allow for a time delay $\Delta t$ in the effect of the wavefunction (having $\left|\psi\left(r,t-\Delta t\right)\right|$ instead of the instantaneous effect $\left|\psi\left(r,t\right)\right|$) to avoid a self-interaction artifact (to be discussed in more detail below). Standard statistical mechanics arguments [21] imply that, in equilibrium at a fictitious temperature $\tau$ (expressed in energy units), the position $r$ will be distributed as $\exp\left(-\left(-\tau\ln\left|\psi\left(r,t\right)\right|^{2}\right)/\tau\right)=\left|\psi\left(r,t\right)\right|^{2}$, as desired. This presupposes that the velocities are initialized (in the distant past) consistently with the fictitious temperature $\tau$, e.g. $\frac{dr_{k}}{dt}\sim\operatorname{Normal}\left(0,\mu_{k}^{-1}\tau\right)$ for $k=1,\ldots,3N$. The velocity degrees of freedom can play the role of thermal bath, so no additional thermal bath is needed. This argument also assumes that the particle equilibrates rapidly relative to the time scale of changes in the wave function magnitude $\left|\psi\left(r,t\right)\right|$. This equilibration behavior is implied by Boltzmann’s classic “H-theorem”, although it is difficult to formulate simple general conditions guaranteeing a given rate of convergence (for a recent account of these types of results, see [22]). In fact, in the sequel we assume that such fast equilibration behavior is guaranteed only for microscopic systems. A simple dimensional argument indicates that the equilibration time should scale with $\ell/\sqrt{\mu^{-1}\tau}$, where $\ell$ is of the order of magnitude of the linear spatial extent probability distribution of a particle’s position.
Note that all quantities labelled here as “fictitious” are not tied to any real observable properties of the system: $\mu$ is not necessarily related to the actual particles’ masses and $\tau$ may not be the real temperature. In fact, these quantities enter all equations as the ratio $\mu^{-1}\tau$, which we leave as an adjustable parameter in the theory that controls the equilibration time. It is interesting that the logarithm in Equation (2) coincidentally accomplishes two tasks: separating out the magnitude from the phase (since $\ln\left|\psi\left(r,t\right)\right|^{2}=2\operatorname{Re}\ln\psi\left(r,t\right))$ and placing the magnitude on a logarithmic scale, as needed to cancel out the Boltzmann factor.
[…]
Figure 1: (Color online) Microscopic vs. macroscopic behaviors. Upper left: For a spatially localized wavefunction, the Bohmian particle (represented by a grey spheres trail) can access all regions where the wavefunction is nonnegligible (here represented by fictitious potential surface proportional to $\ln|\psi(r,t)|^{2}$). Middle: As a superposition of states starts to involve spatially well-separated states, the particle remains trapped in one of the “lobes” of the wavefunction. Lower right: Loss of ergodicity causes wavefunction localization around one of the macroscopically distinct outcomes
Traditional Bohm-de Broglie dynamics guarantees a strong equivariance property: If an ensemble of Bohmian particles is distributed according the $\left|\psi\left(r,t_{0}\right)\right|^{2}$ at some time $t_{0}$, it will remain, by construction, distributed according to $\left|\psi\left(r,t\right)\right|^{2}$ for all $t>t_{0}$. However, we deviate from standard Bohm-de Broglie dynamics for two reasons. First, the particles tend to not get trapped very well under the usual Bohm-de Broglie dynamics: As the particle approaches a region of low probability density, it typically speeds up (which specifically ensures that the particle has low probability of being found there). As a result, the particle can easily escape any isolated lobe in the wavefunction. Second, the Bohmian particle does not thermalize very quickly. For instance, in a wave packet, the particle typically follows along smoothly while remaining at the same location relative to the center of the wave packet, instead of visiting all positions of high probability. Our alternative dynamics avoids both of these problems. The equivariance property under Bohm-de Broglie dynamics is here replaced by a convergence towards $\left|\psi\left(r,t\right)\right|^{2}$ implied by Boltzmann’s “H-theorem” [22]. When the relaxation time associated with this process become very long, the particle’s distribution could differ from $\left|\psi\left(r,t\right)\right|^{2}$ and this possibility will in fact turn out to be essential to enable wavefunction collapse in macroscopic systems in our formalism.
We are now ready to show that, under conditions ensuring rapid thermalization (i.e. when the particle’s relaxation time is sufficiently short), our proposal essentially reduces to the conventional Schrödinger equation. Indeed, computing the expected value of the last term of (1) with respect to the random position $r$ with instantaneous probability density $\left|\psi\left(r,t\right)\right|^{2}$, we have:
$\int_{r\in\mathbb{R}^{3N}}\kappa\bar{\phi}_{r}\left(x\right)\frac{1}{\left|\psi\left(r,t\right)\right|^{2}}\left|\psi\left(r,t\right)\right|^{2}\psi\left(x,t\right)\penalty 10000\ dr$
$=\kappa\frac{1}{\#\Pi}\sum_{\pi\in\Pi}\int_{r\in\mathbb{R}^{3N}}\phi\left(x-\pi r\right)dr\penalty 10000\ \psi\left(x,t\right)=\kappa\psi\left(x,t\right)$ (3)
since $\int\phi\left(x-\pi r\right)dr=1$ by construction. As a result, on average, the last term cancels out the second term of (1) and we recover Schrödinger’s equation.
In contrast, as illustrated in Figure 1, if the wavefunction develops spatially well-separated lobes and the particle is trapped in only one of the them (i.e. when the particle’s global relaxation time becomes very long), then the last term remains essentially zero for the lobes that are no longer visited and, there, only the decaying term $\kappa\psi\left(x,t\right)$ is active, resulting in an exponential decay of the wavefunction. Within the lobe trapping the particle, the localization term of (1) is larger than average, since the particle visit points there in more often than would be indicated by the wavefunction magnitude. The end result is that the localization terms there exceeds the decaying term on average and the wavefunction gradually concentrates onto that lobe at the expense of the others. The loss of ergodicity of the Bohmian particle is the source of the wavefunction collapse within this model. This helps explain single outcomes in macroscopic measurements, since they arguably lead to wavefunctions with macroscopically disconnected lobes, such as a dial (or “pointer”) being in one place or another or a photon being emitted in different directions, etc. Appendix B extends this reasoning to the measurement of a subsystem.
Similarly, another mechanism that could cause a loss of ergodicity is the process of copying a microscopic state onto a large number of particles (within the limits of the no-cloning theorem). This operation typically results in increasingly deeper basins in the fictitious potential $\tau\ln\left|\psi\left(r,t\right)\right|^{2}$ that increase the likelihood of the particle being trapped. To see this, consider a particle in a superposition of two states $\alpha\left|1\right\rangle+\beta\left|2\right\rangle$ (with $\left|\alpha\right|^{2}+\left|\beta\right|^{2}=1$) that is starting to interact with a measurement apparatus made of N1N-1 distinguishable particles, initially in a “no measurement” state $\left|0\right\rangle\otimes\cdots\otimes\left|0\right\rangle$. After a suitable unitary evolution, the joint system transitions to the state $\alpha\left|1\right\rangle\otimes\left|1\right\rangle\otimes\cdots\otimes\left|1\right\rangle+\beta\left|2\right\rangle\otimes\left|2\right\rangle\otimes\cdots\otimes\left|2\right\rangle$. Now assume that, for particle $i$, the states $\left|1\right\rangle$ and $\left|2\right\rangle$ are, respectively, associated with single-particle wavefunctions in position representation $\psi_{1}\left(x_{i}\right)=\varphi\left(x_{i}-a\right)$ and $\psi_{2}\left(x_{i}\right)=\varphi\left(x_{i}+a\right)$ for some macroscopic value $a\in\mathbb{R}^{3}$ and centrosymmetric unimodal function $\varphi\left(x_{i}\right)$ (with $x_{i}\in\mathbb{R}^{3}$). The overall wavefunction is then
$\psi\left(x\right)=\alpha\prod\nolimits_{i=1}^{N}\varphi\left(x_{i}-a\right)+\beta\prod\nolimits_{i=1}^{N}\varphi\left(x_{i}+a\right).$
To estimate the escape rate out of one well, we calculate the fictitious energy barrier $\Delta\epsilon$, defined as the difference between the particle’s fictitious energy at the saddle point at $x_{i}=0$ (for all $i$) and at the bottom of one of the wells (say, at $x_{i}=a$, for all $i$): $\Delta\epsilon\equiv C+N2\tau\ln\left(\left|\varphi\left(0\right)\right|/\left|\varphi\left(a\right)\right|\right)$, where we assumed $\varphi\left(2a\right)\approx 0$ and $C$ is a constant independent of $N$. Since the barrier increases linearly with $N$, the transition rate between the two wells is decreasing exponentially with the number $N$ of particles involved in the measurement.
Figure 2 illustrates our proposed time evolution of the wavefunction in a classic setting: the double-slit experiment. The simulation shows both the initial wave-like interference behavior and the eventual collapse to a more specific outcome as the wave packet heads towards a detector. The Supplementary materials include an animation of the same numerical experiment, as well as the simulation code itself. Appendix A also includes a treatment of a standard EPR-type experiment [23].
[…]
3 Discussion
The proposed theory improves upon the Bohm-de Broglie pilot-wave approach along a few dimensions. First, it avoids the awkwardly distinct treatment of (i) the physical system being described by the wavefunction and (ii) the observer only perceiving the position of the Bohmian particle. Having a feedback mechanism through which the particle affects the wavefunction solves this problem and also leads to a model of wavefunction collapse. In our approach, both the wavefunction and the particle jointly describe the behavior of both the physical system and the observer. The definite state of the particle is what drives the wavefunction localization during the measurement process. A conceptually appealing feature of the theory is that avoids the often criticized exponential growth of alternative realities in which the Bohmian particle does not reside and that seems entirely superfluous to describe the actual state of the world. In our approach, the wavefunction associated with these alternative realities simply die out exponentially fast.
The proposed theory also improves some aspects of objective collapse theories, in which the position of the collapse is simply picked at random based on the wavefunction and a random noise field. On a conceptual level, our approach provides a rationale for the origin of this randomness: The random location of the collapse is simply determined by the random location of the particle. The type of collapse we propose may also be more amenable to an eventual relativistic treatment, as we shall discuss further below. On a more empirical level, a problem with objective collapse theories is the abruptness of the collapse process, which is associated with a high-frequency signature which has yet to be found experimentally [3]. In our approach, the collapse is a smoother process more consistent with experimental tests.
The idea that the wavefunction is attracted to particles can be found in earlier work [18, 24], but our approach differs in substantial ways. Unlike [24], we show that the two-way coupling leads to Born’s rule. Unlike [18], we consider collapse towards a point in $\mathbb{R}^{3N}$ instead of a towards a single-particle density, which helps maintain EPR-type correlations between distant measurements. Also, to ensure particle trapping, our particle dynamics differs from those of both [24] and [18].
The proposed theory is compatible with the standard decoherence picture (see Appendix D for a density matrix formulation): it does not affect the basic idea that interactions with the environment causes a decay of the off-diagonal terms of a subsystem’s density matrix. But once the resulting mixture of outcomes involves macroscopically distinct states, our equation of motion predicts localization towards one of these macroscopic states. Hence, it provides an objective “selection rule” that is absent under decoherence alone and avoids the need for a many-world interpretation. The decoherence concept makes it unnecessary for the wave function to collapse rapidly in our framework. The conversion from a superposition of states to a statistical mixture takes place over the same time span under our theory as under decoherence alone. The exponential decay of all but one branch of the statistical mixture, however, can take much longer: any decay rate faster than human perception (i.e., milliseconds) would be sufficient to give human observers the experience of a single measurement outcome.
The proposed theory does exhibit some undesirable features (see also Appendix C). Most obviously, the proposed modified Schrödinger equation is slightly nonlinear, whereas linearity of quantum mechanics has been verified experimentally to a great precision [25]. This places a tight upper bound on the parameter $\kappa$, thus also placing a bound on the speed at which wavefunction collapse can occur. The fact that the nonlinearity must be small mitigates other known issues associated with nonlinearity [26]. The most troubling of these issues is the possibility of faster-than-light communication. However, the objective collapse literature [27] already offers a natural solution to this possible problem that is relevant within our approach as well. Theorem 1 in [27] provides conditions under which faster-than-light communication is precluded by a suitable combination of nonlinearity with randomness. The key condition is that nonlinearity and randomness cancel out on average, so that the expected time evolution of the density matrix remains a linear operator, where the expectation is taken with respect to a random collapse field. This condition is exactly satisfied in objective collapse theories, in large part because the value of the random collapse field at any two different times are completely statistically independent. In our formalism, the role of the effective random collapse field is played by a smoothed delta function centered at the Bohmian particle’s location $r\left(t\right)$. Over a short time scale, such field unfortunately does not exhibit statistical independence. Yet, over a time scale larger than the particle’s relaxation time, the current particle’s position becomes statistically independent from the particle’s position at a given time in the past. It is known that measure-preserving dynamical systems (such as our classical particle at $r\left(t\right)$ in a potential $\tau\ln\left|\psi\left(r,t\right)\right|^{2}$), typically lead to mixing processes [28]. Hence, for sufficiently large time lags, our effective random field behave analogously to that of objective collapse theories. In addition to this observation, the expected time evolution of the wavefunction under our dynamics either reduces to Schrödinger’s equation (when ergodicity holds), as shown in Equation (3), or reduces to wavefunction collapse (when ergodicity fails). This implies that the density matrix evolution remains linear and thus preclude faster-than light communication in that time regime. Within our theory, detecting faster-than light communication would entail tracking changes in the wave function over very short time scales. At the same time, if the parameter $\kappa$ is small ($\kappa\ll\sqrt{\mu^{-1}\tau}/\ell$), wavefunction evolution over short times closely follows Schrödinger’s equation and the window through which faster-than-light communication might be detected is expected to be very narrow.
While our current treatment is non-relativistic, many aspects of our framework are amenable to a relativistic extension. The collapse phenomenon involves local operations: the decay term $\kappa\psi\left(x,t\right)$ is purely local while the localization terms only involves instantaneous effects over short distances (as governed by the function $\phi_{r}\left(x\right)$) and ensuring that this effect propagates at a finite speed entails no significant changes in the theory. Since relativistic versions of pilot-wave [29, 24] and objective collapse theories [30] have been developed with some success, similar ideas could carry over in the present setup. However, the presence of a global hidden variable (the particle’s position $r$) could clash with relativity, but only over time scales shorter than the particle’s relaxation time.
In our theory, the time delay $\Delta t$ should satisfy $\Delta t\gg\sigma/\sqrt{\tau/\mu}$ (with scalar mass $\mu$ for simplicity) to avoid a strong instantaneous self-interaction. In this fashion, the wavefunction localization only starts to have an effect on the particle when it has already moved by $\Delta t$ times its average velocity $\sqrt{\tau/\mu}$. If this distance exceeds $\sigma$, the particle will not experience the “indentation” it has just created in the wavefunction. Absent of this delay, the particle would have a tendency to gradually lose kinetic energy. This is not the only way to address the self-interaction problem: one could add a negative drag term in the equation of motion for $r$ instead, or place the particle in weak interaction with an outside reservoir that constantly replenishes the lost energy to maintain an average kinetic energy $\tau$. We favor the time delay approach for its simplicity, but it would be difficult to discern these alternatives experimentally.
4 Conclusion
The key new aspects of the theory include: (i) a mutual feedback between the Bohmian particle and the wavefunction; (ii) a different proposed equation of motion for the Bohmian particle; (iii) a mechanism based on the loss of ergodicity (or trapping) of the Bohmian particle to explain to wavefunction collapse phenomena; (iv) the idea that the collapse mechanism can actually be gradual and local. An interesting feature of our approach is that it explores slight relaxations of some properties that are strictly imposed in other theories. The equivariance of the particle’s distribution found in Bohmian-de Broglie pilot wave theories is replaced by fast convergence towards an equilibrium distribution in microscopic systems while macroscopic systems can experience loss of ergodicity. The strict prohibition of faster-than-light communication in standard quantum mechanics and objective collapse theories is relaxed so that superluminal features are allowed, but constrained to a small and difficult-to-access time regime window. Finally, the strict linearity of quantum mechanics is slightly relaxed and this fact could enable tests of the proposed theory.
Perhaps the most direct test would involve generating particles in a superposition of two spatially-separated wavepackets-like states that are later recombined, and then analyzing how contrast of the resulting interference patterns depend on the spatial separation of the initial wavepackets. Very sensitive experiments of this kind have been performed [31], whose primary finding was to demonstrate that interference contrast can survive even for wavepacket separated by about half a meter for about a second. However, as a ancillary benefit, such experiments also revealed a strong dependence of that contrast on the wavepacket separation (see Figure 4b in [31]). This observation is consistent with our proposed mechanism but not entirely conclusive, as it could also arise for other reasons, such as decoherence becoming more pronounced for more distant wavepackets. Hence, a rigorous test would involve comparing processes that are expected to undergo a similar amount of decoherence but that differ in the length and/or time scale involved. Such confirmation would shed light on some of the conceptual foundation of quantum mechanics and could inform us regarding the feasibility of large-scale quantum computers.
Appendix A EPR-type experiments
It is instructive to see how the theory handles EPR-type experiments [23]. In this case, the long-distance correlations in the wavefunctions simply induce similar long-distance correlations between different component of the $3N$-dimensional particle position. As each observer of one particle of the EPR pair makes a measurement, they each create disjoint lobes in their wavefunction. Specifically, consider an EPR pair is in a state $\left(\left|H\right\rangle\otimes\left|H\right\rangle+\left|V\right\rangle\otimes\left|V\right\rangle\right)/\sqrt{2}$ and one observers measures the first particle in the basis $\left|H\right\rangle,\left|V\right\rangle$ (e.g., horizontally or vertically polarized photon), while the other observer measures the second particle in the basis $\left|P\right\rangle=\cos\left(\theta\right)\left|H\right\rangle+\sin\left(\theta\right)\left|V\right\rangle$, $\left|M\right\rangle=-\sin\left(\theta\right)\left|H\right\rangle+\cos\left(\theta\right)\left|V\right\rangle$. Expressed in the observers’ basis the pair’s state is
$\frac{\cos\left(\theta\right)}{\sqrt{2}}\left|H\right\rangle\otimes\left|P\right\rangle-\frac{\sin\left(\theta\right)}{\sqrt{2}}\left|H\right\rangle\otimes\left|M\right\rangle+$
$+\frac{\sin\left(\theta\right)}{\sqrt{2}}\left|V\right\rangle\otimes\left|P\right\rangle+\frac{\cos\left(\theta\right)}{\sqrt{2}}\left|V\right\rangle\otimes\left|M\right\rangle.$
Now, we consider a simple measurement process that destructively copies the state of the pair onto two macroscopic pointers (initially in some neutral state $\left|\psi_{10}\right\rangle\otimes\left|\psi_{20}\right\rangle$ ). Specifically, the measurement process maps, for particle $p=1,2$, the state $\left|s\right\rangle\otimes\left|\psi_{p0}\right\rangle$ (for $s=H,V,P,M$) onto the state $\left|0\right\rangle\otimes\left|\psi_{ps}\right\rangle$, where $\left|0\right\rangle$ is some arbitrary chosen neutral state (it could be any one of $H,V,P,M$ — as long as it is always the same state regardless of the measurement outcome). (As an alternative, we could also consider measurement that maps $\left|s\right\rangle\otimes\left|\psi_{0}\right\rangle$ onto the state $\left|s\right\rangle\otimes\left|\psi_{s}\right\rangle$, at the expense of necessitating an additional tracing out operation.)
After the measurement, the system and the pointers are in the state
$\left|0\right\rangle\otimes\left|0\right\rangle\otimes\left(\frac{\cos\left(\theta\right)}{\sqrt{2}}\left|\psi_{1H}\right\rangle\otimes\left|\psi_{2P}\right\rangle-\frac{\sin\left(\theta\right)}{\sqrt{2}}\left|\psi_{1H}\right\rangle\otimes\left|\psi_{2M}\right\rangle+\right$.
$\left.+\frac{\sin\left(\theta\right)}{\sqrt{2}}\left|\psi_{1V}\right\rangle\otimes\left|\psi_{2P}\right\rangle+\frac{\cos\left(\theta\right)}{\sqrt{2}}\left|\psi_{1V}\right\rangle\otimes\left|\psi_{2M}\right\rangle\right)$
whose probability density, expressed in the position basis for the pointers, is:
$\frac{\cos^{2}\left(\theta\right)}{2}\left|\psi_{1H}\left(x_{1}\right)\right|^{2}\left|\psi_{2P}\left(x_{2}\right)\right|^{2}$
$+\frac{\sin^{2}\left(\theta\right)}{2}\left|\psi_{1H}\left(x_{1}\right)\right|^{2}\left|\psi_{2M}\left(x_{2}\right)\right|^{2}+$
$+\frac{\sin^{2}\left(\theta\right)}{2}\left|\psi_{1V}\left(x_{1}\right)\right|^{2}\left|\psi_{2P}\left(x_{2}\right)\right|^{2}$
$+\frac{\cos^{2}\left(\theta\right)}{2}\left|\psi_{1V}\left(x_{1}\right)\right|^{2}\left|\psi_{2M}\left(x_{2}\right)\right|^{2},$
where we have assumed negligible overlap between the wavefunctions associated with different pointer positions ($\psi_{ps}\left(x\right)\psi_{ps^{\prime}}\left(x\right)\approx 0$ for $s\not=s^{\prime}$). As these 4 lobes of probability have little spatial overlap, the Bohmian particle would get trapped in one of them, with a probability equal to that particle being found in that lobe at the time they became disconnected. The end results are measurements with outcomes $\left(H,P\right),$ $\left(H,M\right),$ $\left(V,P\right),$ $\left(V,M\right)$ with probability $\cos^{2}\left(\theta\right)/2,\sin^{2}\left(\theta\right)/2,\sin^{2}\left(\theta\right)/2,\cos^{2}\left(\theta\right)/2$, respectively, in agreement with Born rule and the known outcome of EPR-type experiments. Note that there are no contradictions with Bell inequalities because the hidden variables $r$ are not local.
Appendix B Subsystem Measurements
It is instructive to consider how the theory handles the fact that one typically seeks to measure the state of only some subsystem and not of the universe as a whole. In Figure 3, the $x_{1}$ coordinate describes the degree of freedom to be measured, while the $x_{2}$ coordinate represents the rest of the universe. The degrees of freedom that are being measured are simply those along which the wavefunction develops disconnected lobes. The remaining degrees of freedom remain fully connected and the particle can continue to reach its equilibrium distribution along those dimensions. The net effect is that the remaining degrees of freedom still closely follow Schrödinger’s equation and only the degrees of freedom that are being measured experience significant deviations from unitary evolution. A similar argument can be made for “weak” measurements.
[…]
Figure 3: (Color online) Schematic representation of the measurement of the position $x_{1}$ of one particle, while other degrees of freedom (represented by $x_{2}$) are left unaffected. Connectivity is lost only along the $x_{1}$ dimension. (Particles are considered distinguishable for simplicity.)
Appendix C Potential Objections
Just like both pilot-wave and objective collapse theories, our proposal appears to clash with conventional quantum mechanics by introducing a preferred basis: position eigenstates have a special status compared to eigenstate of any other observables. This does not present a significant problem, to the extent that a measurement apparatus typically maps different values of any observables of interest onto macroscopic states with different spatial distribution. The latter would initiate the selection of one outcome based on position and the microscopic system of interest would inherit this choice through its entanglement with the measurement outcome. This argument holds even if the microscopic observable of interest is not position. Furthermore, it can be argued that there is some experimental support to the idea that superpositions of states involving different spatial distributions are more difficult to obtain. Indeed, superconducting qubits can now routinely exist in macroscopic (micrometer-size) superpositions of different spin or momentum states [32]. In contrast, placing a macroscopic mechanical oscillator in a superposition of states that differ my $\sim 10^{-18}\,$m is considered a remarkable accomplishment [33]. Nevertheless, it is possible to modify the theory to allow for collapse onto a preferred basis other than position — see Discrete States section below.
An unusual feature of our approach is the lack of time-reversal symmetry of our wave equation (1), due to the presence of terms without an imaginary prefactor in the first derivative expression. A side-effect of this would be to cause the quantum entropy of a subsystem to increase more slowly than under the standard Schrödinger equation because the wavefunction has a tendency to localize. This does not necessarily entail a violation of the second Law of thermodynamics because it merely converts quantum entropy (due to the subsystem not being in a pure state) into a classical entropy (reflecting the observer’s lack of information regarding the actual state of the subsystem). From a macroscopic thermodynamic point of view, either interpretation equally contributes to the macroscopic entropy.
Appendix D Density Matrix Formulation
Our equations of motion can readily be expressed in terms of the density matrix $\rho$:
$\frac{d\rho}{dt}$ $=\frac{1}{\mathbf{i}\hbar}\left[H,\rho\right]-2\kappa\rho+\kappa\frac{\left[\left|\bar{\phi}_{r}\right\rangle\left\langle\bar{\phi}_{r}\right|,\rho\right]_{+}}{\left\langle r\right|\rho\left|r\right\rangle}$
[…]
This density matrix formulation is very helpful to trace out non-positional degrees of freedom (such as spins) and cleanly obtain an equation of motions for the positional degrees of freedom only.
In conventional quantum mechanics, one can freely “trace out” the density matrix to obtain a separate equation of motion for any subsystem, provided there is no coupling between the subsystem and the environment in the Hamiltonian $H$. This works even if there is entanglement between the subsystem and the environment. In our approach, there is also no problem associated with tracing out the density under these conditions, but the particle position $r$ cannot always be separated out so easily, a problem which does not occur when tracing out over non positional degrees of freedom. If there is entanglement between the subsystem and the environment, the trajectory of the particle within the subsystem is still affected by the environment (through the $\left\langle r\right|\rho\left|r\right\rangle$ term), even if there is no coupling through the Hamiltonian. An independent equation of motion for subsystem can only obtained if the overall density matrix factors as a product of the density matrix of the system and that of the environment. This is an interesting distinction with conventional quantum mechanics.
Appendix E Discrete States
A potentially useful extension of our theory would be to include systems with discrete states where there is no natural definition of position and momentum. This would allow for wavefunction collapse in spin systems or fermionic quantum fields, for instance, without resorting to the argument that the collapse only occurs when such systems are coupled with a macroscopic position-based pointer.
In some sense, this task is far easier to do with our approach than with standard pilot-wave theory, because our equation of motion for the particle position $r$ does not rely on the definition of a momentum operator. A natural discrete formulation of our modified Schrödinger equation would be:
$\frac{d\psi\left(t\right)}{dt}=\frac{1}{\mathbf{i}\hbar}H\psi\left(t\right)-\kappa\psi\left(t\right)+\kappa\frac{\delta_{r}}{\left|\psi_{r\left(t\right)}\left(t\right)\right|^{2}}\psi\left(t\right)$
where the wave function $\psi\left(t\right)$ is represented by a vector with elements $\psi_{r}\left(t\right)$, while $r\left(t\right)$ represents the current discrete state of the Bohmian particle at time $t$. The localization function $\bar{\phi}_{r}$ just becomes a Kronecker delta $\delta_{r}$. It is not clear how to obtain a deterministic equation of motion for $r\left(t\right)$, but one could postulate a stochastic evolution equation (paralleling [34]) that specifies the probability of transition $P_{r^{\prime}r}$ from a state $r$ to another state $r^{\prime}$ per unit time. The transition probability matrix $P_{r^{\prime}r}$ must satisfy the condition that it admits an attractive equilibrium distribution of $r$ that matches $\left|\psi_{r}\right|^{2}$. The theory of Monte Carlo sampling provides many convenient valid choices, for instance:
$P_{r^{\prime}r}\propto\left(1+\frac{\left|\psi_{r}\right|^{2}}{\left|\psi_{r^{\prime}}\right|^{2}}\right)^{-1}K_{r^{\prime}r}$
[…]